4. (5 points) Data from an experiment yielded the following graph. What is the resistance of the resistor? I (A) 0.025 Ohm V (V) Explain your reasoning
Answer #1

According to Ohms law,

I is proportional to V

and the proportioanlity constant is the inverse of Resistance, hence the equation becoms

I=V/R

V=IR

according to the equation, the slope of the graph would be

m=1/R

R= 1/m = 1/0.025

R=40 ohms